Find a polynomial of least degree with integer coefficients that has the given zeros. Write your answer in both factored form and general form.
X= 0,-4,1

Answer by ikleyn(52851)

.

In factored form, the polynomial is the product of the three linear binomials associated with the given roots


    p(x) = (x-0)*(x-(-4)*(x-1) = x*(x+4)*(x-1).


You can do the rest.